f(x)=√(x-3)(2+5x)
Im not sure how to solve it as well as to finding the domain

Answer by stanbon(75887)

f(x)=√(x-3)(2+5x)
Im not sure how to solve it as well as to finding the domain
---
1st: You cannot "solve it".
---
2nd: Domain:
(x-3)(2+5x) must not be negative:
Solve (x-3)(2+5x) >= 0
---
The left side is zero when x = -2/5 or when x = 3
----
Draw a number line and plot x = (-2/5) and x = 3
-----
Test for solution areas on the number line:
(x-3)(2+5x) > 0
If x = -10 ; -*- > 0 ; true, so solution in (-oo,-5/2)
If x = 0 ; -*+ > 0 ; false
If x = 10 ; +*+ > 0 ; true, so solutions in (3,+oo)
-----
Domain: (-oo,-5/2]U[3,+oo)
===============================
cheers,
Stan H.